OpenStudy (anonymous):

A quadratic equation is shown below: x^2 + 18x + 76 = 0 Which of the following is the first correct step to write the above equation in the form (x - p)^2 = q, where p and q are integers?

OpenStudy (anonymous):

Add 9 to both sides of the equation Add 5 to both sides of the equation Subtract 5 from both sides of the equation Subtract 9 from both sides of the equation

OpenStudy (anonymous):

\[x^2 + 18x + 76 = 0\]\[x^2 + 18x + 76\color{red}{+5} = 0\color{red}{+5} \]and then\[x^2 + 18x + 81 =5\]\[(x+9)^2 = 5\]\[x+9 = ±~5\]\[x = ±~5-9\]\[x =14,~~~~~~or~~~~~x=-4\]

OpenStudy (zzr0ck3r):

I think there was a typo in the question, it says (x-9)^2

OpenStudy (zzr0ck3r):

err (x-p) nm....

OpenStudy (zzr0ck3r):

dislexia....

OpenStudy (secret-ninja):

+5 to both sides

OpenStudy (anonymous):

Its add 5 to both sides, I took that test too.....